One side of a triangle is half the longest side. The third side is 12 feet less than the longest side. The perimeter is 53 feet. Find all three sides!
Found 2 solutions by lwsshak3, webladyusa:
Answer by lwsshak3(11628)

One side of a triangle is half the longest side. The third side is 13 feet less than the longest side. The perimeter is 62 feet. Find all three sides
**
let x=longest side
x/2=one of the sides
x-13=3rd side
..
x+x/2+x-13=62
2.5x=75
x=30
..
Ans:
longest side=30 ft
one other side=15 ft
3rd side=17ft

Answer by webladyusa(1)

One side of a triangle is half the longest side. The third side is 12 feet less than the longest side. The perimeter is 53 feet. Find all three sides!
x+2x+2x-12=53
5x=53+12
5x=65
x=13
2x=26
2x-12=14
Longest side=26 ft
One other side=13 ft
3rd side=14ft